OpenStudy (anonymous):

Given that f(x) = 5x – 10 and g(x) = x + 3, solve for f(g(x)) when x = –1. –30 0 10 12

OpenStudy (freckles):

find g(-1) first

OpenStudy (freckles):

notice g(x)=x+3 enter in -1 for x

OpenStudy (jdoe0001):

actaully is hmm -1 well

OpenStudy (freckles):

your work isn't meaningless though it is good example

OpenStudy (jdoe0001):

\(\bf f(x)=5x-10\qquad g(x)=x+3 \\ \quad \\ f(\quad {\color{brown}{ g(x)}} \quad )=5{\color{brown}{ g(x)}}-10\implies f(\quad {\color{brown}{ g(x)}} \quad )=5{\color{brown}{ (x+3)}}-10 \\ \quad \\ f(\quad {\color{brown}{ g({\color{blue}{ -1}})}} \quad )=5{\color{brown}{ g({\color{blue}{ -1}})}}-10\implies f(\quad {\color{brown}{ g({\color{blue}{ -1}})}} \quad )=5{\color{brown}{ ({\color{blue}{ -1}}+3)}}-10\)

OpenStudy (freckles):

ok nevermind

OpenStudy (jdoe0001):

:)

OpenStudy (anonymous):

12?